Solve by completing the square.
3x^2-3x-6=0

Answer by edjones(8007)

3x^2-3x-6=0
x^2-x-2=0 divide each side by 3
x^2-x =2
x^2-x+ 1/4 = 9/4 divide the 2nd term (-1) by 2 and square the result.
x- 1/2 = +-3/2 sqrt of each side.
x=1/2 + 3/2 =2, x=1/2- 3/2 =-1
.
Ed